OpenStudy (anonymous):

Test the convergence of the series 3^n+4^n/4^n+5^n using limiting form method...the Range of n starts from 1

OpenStudy (anonymous):

divide all by 5^n

OpenStudy (anonymous):

every single term

OpenStudy (anonymous):

By limiting form method we need vn which is larger than un the given series..so how to find vn

OpenStudy (anonymous):

@UncleRhaukus

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

@Jemurray3 @.Sam. @Shane_B could you help me please

OpenStudy (dape):

Always use parenthesis! If you meant \[ \sum_{n=1}^{\infty}{3^n+\frac{4^n}{4^n}+5^n} \] the series diverges, if you meant \[ \sum_{n=1}^{\infty}{\frac{3^n+4^n}{4^n+5^n}} \] You can use the ratio test to test it's convergence.

OpenStudy (anonymous):

i meant the second one @dape ..and how to do this ratio test

OpenStudy (dape):

Each term in the series is given by \[ a_n = \frac{3^n+4^n}{4^n+5^n} =\frac{\frac{1}{5^n}(3^n+4^n)}{\frac{1}{5^n}(4^n+5^n)} =\frac{\frac{3^n}{5^n}+\frac{4^n}{5^n}}{\frac{4^n}{5^n}+\frac{5^n}{5^n}} = \frac{(\frac{3}{5})^n+(\frac{4}{5})^n}{(\frac{4}{5})^n+1} \] If the sum of all these terms is to converge, the terms much get smaller and smaller in the limit as n get bigger and bigger. In other words we must have \[ \lim_{n\rightarrow\infty}\frac{a_{n+1}}{a_n} < 1 \] Try to check if the above statement is true.

OpenStudy (anonymous):

so the series diverges?

OpenStudy (dape):

Nope, it converges, as you will se if you evaluate the limit.

OpenStudy (anonymous):

actually i did it like this i made \[4^n\] as \[(3+1)^n\] and took out \[3^n\] as common and did the same with \[5^n\] as \[(4+1)^n\] and finally made \[({3/4})^n\] as Vn and found the limit of Un/Vn where Un is \[({3/4})^n *( 1+({1+1/3})^n/1+({1+1/4})^n)\]

OpenStudy (anonymous):

and then applied the limits for Un/Vn (n--->infinity)and got it as 1.

OpenStudy (dape):

You can't take out \(3^n\) from \((3+1)^n\) without getting other nasty terms. The correct limit I got to 0.8, which means it converges.